Ashley works as a tutor for
$8
an hour and as a waitress for
$14
an hour. This month, she worked a combined total of
87
hours at her two jobs.

Answer by addingup(3677)

tutor T
waitress W
----------------------
T + W = 87; T = 87 - W
8T + 14W = ? you need this information. How much did she make this month?